OpenStudy (anonymous):

write an equation of the circle whose diameter line ab has endpoints a(-4,2) and b(4,-4) full answer show all work poleaseeeee

OpenStudy (amistre64):

-4+4 = 0/2 = 0 2-4 = -2/2 = -1; the center is at (0,-1) our radius is sqrt((0-4)^2+{-1--4}^2) = sqrt(16+9) = 5 x^2 + (y+1)^2 =25 if i see it right

OpenStudy (anonymous):

in a diagram quadrilateral STAR is a rhombus with diagonals line SA and line TR intersecting at E. ST= 3x+30 ,SR= 8x-5, SE=3z,TE=5z+5, AE=4z-8, m<RTA=5y-2 and m<TAS=9y+8 find SR,RT, and M<TAS show all work thankss

OpenStudy (amistre64):

sorry, i can t make sense of that

OpenStudy (anonymous):

okay thank you anyway